C 1 7. [8] AABC and ADEB are both right triangles. AC = 3 AB = 4 Also, D is the midpoint of AB. Find DE E A B D
Answer #1

Solution:

In triangle ABC

AC=3

AB=4

\\BC=\sqrt{3^2+4^2}=\sqrt{9+16}=\sqrt{25}=5\\BC=5

BC=5

In triangle BDE

BD=2 (D is the midpoint of AB )

In both triangle are similar

\\\frac{BC}{BD}=\frac{AC}{DE}\\\\\frac{5}{2}=\frac{3}{DE}\\\\DE=\frac{2}{5}\cdot 3=\frac{6}{5}=1.2\\\\DE=1.2